OpenStudy (anonymous):

holly wants to save money for an emergency. holly invests $1200 in an account that pays an interest rate of 4%. how many years will it take for the account to reach $15,400? round your answer to the nearest hundredth

OpenStudy (anonymous):

It's 4% per month right. So SI = ( PRT ) /100 15400 -1200 = 1200 * 4 * T /100 14200 = 48 T T = 14200/48 = 296 months